OpenStudy (anonymous):

Two planes leave from Dallas, Texas. The first plane travels east at a rate of 250 mph. The second plane travels west at a speed of 225 mph. How long will it take for them to be 2850 miles apart? 6 hours 2 hours 8 hours 4 hours

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

http://www.purplemath.com/modules/distance.htm

OpenStudy (raden):

let they are take time t hours, so 250t + 225t = 2850 475t = 2850 t= 2850/475 t = 6 hours